OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to x from the second equation is substituted into the first equation. 4x – 3y = –6 –x – 5y = –9 A.4x – 3(–5y + 9) = –6 B.4(–5y + 9) – 3y = –6 C.4x – 3(5y – 9) = –6 D.4(5y – 9) – 3y = –6

OpenStudy (anonymous):

-x=5y-9 x=-5y+9 4(-5y+9) -3y=-6, B

OpenStudy (anonymous):

-x-5y=-9 Add -5y to get -x=5y-9 Divide both side by -1 x=9-5y Substitute in the x into the first equation to get 4(9-5y)-3y=-6 Which is equivalent to answer B